# mFabric: An Efficient and Scalable Fabric for Mixture-of-Experts Training

### v2

@@ 11. Conclusion
This paper presented mFabric, a novel reconfigurable fabric using commodity hardware for large-scale MoE training. At the core of mFabric is the design and implementation of the regionally reconfigurable high-bandwidth domain based on distributed OCS. Through proof-of-concept prototype and large-scale packet simulations, we show that mFabric delivers distributed training performance comparable to state-of-the-art electrical and optical interconnects while significantly reducing networking cost.
This work does not raise any ethical issues.
@@ Appendix A Production Measurement Details
A.1. Setup
The production fabric. We use a Certified Nvidia DGX SuperPOD platform (nvi, [n. d.]b) with 128 H-series (Hopper architecture) GPUs and 128 ConnectX-7 400Gbps NICs in a production data center. The compute fabric is connected in a rail-optimized topology (rai, [n. d.]). We use NCCL (ncc, [n. d.]) to optimize communications on the DGX platform.
Training configurations. We present the training configuration of different MoE models in Table 3.
A.2. Profiling of MoE Models
@@
A.3. Non-uniform token distribution in trained MoE model
@@
Models Mixtral LLaMA-MoE Qwen-MoE
Size 8$\times$7B 6.7B 14.3B
# of MoE blocks 32 32 24
# experts 8 16 64
EP degree 8 16 16
TP degree 4 1 1
PP degree 4 4 4
Seq. len. 4096 4096 4096
Micro-batch size 8 8 8
Table 3. State-of-the-art MoE training configurations.
@@
Figure 19. Average prediction accuracy of mFabric-Copilot.
@@ B.1. Traffic Demand Prediction
mFabric aims to handle the first all-to-all communication in the forward pass with a predictive approach. By default, the OCS topology for this initial communication is either randomly generated (e.g., for the first all-to-all in the first layer) or remains unchanged from the previously used topology (e.g., the first all-to-all in the second layer). The traffic demand prediction algorithm predicts the conditional probability of the traffic matrix, denoting the conditional probability of a token gated to expert $j$ given that it is gated to the expert $i$ in the last layer. With the conditional probability matrix and the empirical token distribution in the previous layer, we can predict the traffic distribution in the current layer.
Matrix Estimation: For each layer, mFabric estimates the conditional probability matrix with the traffic demand records in recent iterations. Focusing on the recent expert load distributions, we employ a weighted average within a fixed window of traffic records in time series. For each layer, the optimization objective is to minimize the square error between the predicted load distribution and the ground truth (for simplicity, we omit the layer index here):
@@
where $k$ is the window size. The transition matrix $P$ is of size $N\times N$, representing the conditional probability of the current layer’s expert load distribution, given the expert load distribution of the previous layer. $X_{i}$, $Y_{i}$ are normalized expert load distribution vectors of two neighboring layers. Each element in $P$ is constrained to be in the range $[0,1]$, and the sum of each column is constrained to be 1 to ensure that $P$ is a valid probability matrix.
@@
We name this method mFabric-Copilot. Figure 19 compares the prediction accuracy of mFabric-Copilot against the aforementioned methods, i.e., the randomly assigned token distribution (uniform bandwidth allocation), and the unmodified token distribution from the previous layers (unchanged topology) on collected traces from measurements. Top $K$ accuracy measures whether mFabric is able to find the top-k activation-intensive experts. We find that mFabric-Copilot exhibits significantly higher accuracy than other counterparts, which implies that mFabric-Copilot can find the most intensive pairs in all-to-all communication with high probability. Therefore, mFabric-Copilot offers an opportunity to proactively reconfigure the topology for the FP’s first all-to-all in advance.
@@ B.2. Details of Topology Reconfigurations
Figure 20 details the runtime reconfiguration methodology of mFabric. Each MoE layer involves four all-to-all communication phases. As discussed in §5.1, mFabric can deterministically characterize the traffic patterns for the second all-to-all communication in the forward pass (FP) and both all-to-all communications in the backward pass (BP).
Consequently, mFabric conceptually reconfigures the topology twice per MoE layer—once during FP and once during BP. However, for the first all-to-all communication in FP, mFabric cannot fully characterize the traffic matrix in advance due to the absence of runtime information at that stage of the iteration. To tackle this issue, mFabric performs an inaccurate reconfiguration based on partial estimates or reuses the topology from the previous layer. It then recalibrates the topology for the second all-to-all in FP with minor OCS adjustments, ensuring a more accurate configuration for subsequent communication phases.
@@ Appendix C ProtoType Details
As a result, we currently exclude this NIC activation time to calculate the actual training time in mFabric testbed experiments. Due to the limited number of GPUs, we cannot train the full MoE models as shown in Table 3. We only run 7 layers of Mixtral 8$\times$7B, 16 layers of LLaMA-MoE, and 12 layers of Qwen-MoE.
@@ Appendix D Large-Scale Simulation Details
D.1. Simulated MoE Models and Parallelization Strategies.
@@ D.2. Cost Analysis Details
D.3. Different EPS Link Options.
The OCS portion of mFabric requires optical transceivers with pluggable fibers to allow optical switching. For the EPS part of mFabric, especially short-reach rack-scale links between the servers and ToR switches, Direct Attach Copper (DAC) cables or Active Optical Cables (AOC) are more cost-effective alternatives to optical transceivers plus fibers (typically used for long-reach links). We analyze the cost implications of these link options in Figure 24. The results show that replacing the EPS links with DAC or AOC slightly reduce the costs for both fat-tree interconnect and mFabric. Most importantly, the cost effectiveness of mFabric is orthogonal to the choices of EPS links, and maintains significant cost advantages over fat-tree topology. For example, with 400 Gbps DAC cables option in a 4096-GPU cluster, mFabric achieves 2.2$\times$ lower total cost compared to fat-tree topology.
@@ D.4. Training Speed Ups of Mixtral Models with Larger Batch Sizes
We further evaluate mFabric’s performance with larger batch sizes using two Mixtral MoE models (Mixtral 8$\times$7B and Mixtral 8$\times$22B). For each model, we test batch sizes of 32 and 64 across varying network bandwidths (100–800 Gbps). As shown in Figure 25, mFabric consistently outperforms TopoOpt under all configurations. Specifically, mFabric achieves an average speedup of 1.8$\times$ for Mixtral-8x7B with a batch size of 32 and 2.0$\times$ with a batch size of 64, as training becomes more communication-intensive compared to the settings in Figure 12. Furthermore, we observe that as link bandwidth increases, mFabric’s performance gradually approaches that of Fat-tree and Rail-optimized architectures.
@@
Figure 26. [Simulation] Comparison of mFabric with different cluster sizes.
Figure 27. [Simulation] Impact of optical degree $\alpha$ in mFabric
@@ D.5. Scalability
We demonstrate the scalability of mFabric in Figure 26(a). The Mixtral 8$\times$7B model is evaluated at 400 Gbps bandwidth, with the cluster size varying from 128 servers to 4,096 servers, covering up to 32768 GPUs. mFabric demonstrates scalability by fundamentally relaxing the port limits of OCS through the design of several decentralized regionally reconfigurable domains, allowing it to scale similarly to a fat-tree topology. Our results show that mFabric scales effectively with increasing number of GPUs, achieving training throughput comparable to both non-blocking Fat-tree and Rail-optimized topologies in terms of tokens processed per second. We further present the performance-cost comparison in Figure 26(b), which shows that mFabric consistently achieves a superior performance-cost trade-off—approximately 2$\times$ higher performance-per-dollar—compared to Fat-tree and Rail-optimized topologies as the number of GPUs increases. This suggests that mFabric maintains the training cost-effectiveness even as the cluster size grows.
@@ D.6. Impact of Optical Degree
We show the impact of the optical degree on mFabric’s performance in Figure 27. We evaluate the Mixtral 8$\times$22B model on a cluster of 128 servers with 100 Gbps link bandwidth. The optical degree $\alpha$ in mFabric is varied to adjust its connectivity in the OCS. We reduce the bandwidth of each electronic port when increasing their number, to ensure a cost-equivalent comparison. Our findings show that, as the optical degree increases, mFabric further reduces iteration time, as more communication-intensive GPU pairs can be provisioned with dedicated high-bandwidth optical circuits.
D.7. Impact of Reconfiguration Time.
@@
To investigate mFabric’s sensitivity to OCS reconfiguration latency, we evaluate the Mixtral 8$\times$22B model on a cluster of 128 servers with 400 Gbps link bandwidth, varying the reconfiguration latency from 1 µs to 10 s. Figure 28 shows the normalized iteration time. mFabric assumes the use of a millisecond-scale reconfigurable OCS (25 ms) in its current implementation. We observe that further reducing reconfiguration latency does not yield significant performance gains, as the OCS reconfiguration process can already be fully hidden during the computation phase. However, provisioned with microsecond-scale reconfigurable OCS, mFabric can enable fully accurate topology reconfigurations for the first all-to-all communication in the forward pass (FP), resulting in marginal performance improvements in this specific phase. On the other hand, when the reconfiguration latency exceeds 1000 ms, performance degrades obviously, as the OCS reconfiguration may not be hidden and starts to block the training process. As a result, mFabric does not perform well with second-scale reconfigurable OCS systems.
Appendix E Look Ahead
E.1. Deploying mFabric with Co-Packaged Optical I/O Directly Attached to GPUs
mFabric as a network architecture is compatible with emerging physical-layer optical technologies.
Evolution of mFabric. As depicted in Figure 6, the current implementation of the regional OCS fabric in mFabric operates alongside the global scale-out domain using commodity NICs and optical transceivers. This configuration relies on NVSwitch and EPS to relay traffic and maintain global network connectivity. However, advances in co-packaged optical I/O (CPO) (lig, [n. d.]; bro, [n. d.]; aya, [n. d.]) technologies, which enable the integration of computational and optical chips into a single board, open the door to a new design paradigm of GPU-to-GPU direct optical communication. Based on this, we envision mFabric evolving towards a regional OCS architecture capable of directly receiving optical signals from xPUs (  e.g., GPU, TPU, NPU,   etc.), as illustrated in Figure 15. Such an optical switching architecture would enable a long-reach, high-speed (e.g., 4 Tbps or more) regional fabric that sits at the boundary of scale-up and scale-out domains, supporting larger-scale switching fabric compared to state-of-the-art NVL72 copper-based interconnects.
E.2. Compatibility with Other OCS Technologies
mFabric’s regional reconfiguration design is orthogonal to the selection of OCS technologies. We note the emergence of new OCS technologies, such as iPronics (ipr, [n. d.]), which offer microsecond-scale reconfiguration latencies and low per-port costs comparable to MEMS-based OCS. These advancements present a promising opportunity for mFabric to reconfigure the topology for every all-to-all communication phase in MoE training, further improving training performance while maintaining cost efficiency. Furthermore, in the main text, we mainly discuss OCS technologies that can actively switch optical signals at the network core. Other emerging OCS techniques, such as tunable laser at endpoints (e.g., transceivers) with arrayed waveguide grating router (AWGR)-based optical switch that pushes reconfigurability into the network endpoints while leaving the network core to be passive (Ballani et al., 2020; Zhang et al., 2024; Ye et al., 2012) is also compatible with mFabric. For example, assigning dedicated wavelengths to critical expert pairs eliminates bandwidth contention for all-to-all EP communication while organizing ring-based topologies for all-reduced communication of TP. However, these AWGR and endpoint-based OCS are primarily in the research stage and not commercially ready for mass production yet, leaving a longer way to be integrated into today’s computer systems.

### v3

@@ 11. Conclusion
This paper presented MixNet, a novel reconfigurable fabric using commodity hardware for large-scale MoE training. At the core of MixNet is the design and implementation of the regionally reconfigurable high-bandwidth domain based on distributed OCS. Through proof-of-concept prototype and large-scale packet simulations, we show that MixNet delivers distributed training performance comparable to state-of-the-art electrical and optical interconnects while significantly reducing networking cost.
Ethics: This work does not raise any ethical issues.
Acknowledgments
We sincerely thank the anonymous SIGCOMM reviewers and our shepherd, Alex C. Snoeren, for their insightful feedback. We also thank Haiyang Chen, Decang Sun, Jipeng Zhang, and the Polatis team for their support in building the testbed. This work is supported in part by the Hong Kong RGC TRS T41-603/20R, ITC ACCESS, TACC (Xu et al., 2025), EmbedWay research project, Beijing Municipal Science and Technology Project No. Z241100004224023, NSFC Excellent Young Scientists Fund Program (Overseas). Zhizhen Zhong and Kai Chen are the corresponding authors.
Appendices are supporting material that has not been peer-reviewed.
@@ Appendix A Production Measurement Details
A.1. Profiling of MoE Models
@@
A.2. Non-uniform token distribution in trained MoE model
@@
Figure 19. Average prediction accuracy of MixNet-Copilot.
@@ B.1. Traffic Demand Prediction
MixNet aims to handle the first all-to-all communication in the forward pass with a predictive approach. By default, the OCS topology for this initial communication is either randomly generated (e.g., for the first all-to-all in the first layer) or remains unchanged from the previously used topology (e.g., the first all-to-all in the second layer). The traffic demand prediction algorithm predicts the conditional probability of the traffic matrix, denoting the conditional probability of a token gated to expert $j$ given that it is gated to the expert $i$ in the last layer. With the conditional probability matrix and the empirical token distribution in the previous layer, we can predict the traffic distribution in the current layer.
Matrix Estimation: For each layer, MixNet estimates the conditional probability matrix with the traffic demand records in recent iterations. Focusing on the recent expert load distributions, we employ a weighted average within a fixed window of traffic records in time series. For each layer, the optimization objective is to minimize the square error between the predicted load distribution and the ground truth (for simplicity, we omit the layer index here):
@@
where $k$ is the window size. The transition matrix $P$ is of size $N\times N$, representing the conditional probability of the current layer’s expert load distribution, given the expert load distribution of the previous layer. $X_{i}$, $Y_{i}$ are normalized expert load distribution vectors of two neighboring layers. Each element in $P$ is constrained to be in the range [0,1][0,1][ 0 , 1 ], and the sum of each column is constrained to be 1 to ensure that $P$ is a valid probability matrix.
@@
We name this method MixNet-Copilot. Figure 19 compares the prediction accuracy of MixNet-Copilot against the aforementioned methods, i.e., the randomly assigned token distribution (uniform bandwidth allocation), and the unmodified token distribution from the previous layers (unchanged topology) on collected traces from measurements. Top $K$ accuracy measures whether MixNet is able to find the top-k activation-intensive experts. We find that MixNet-Copilot exhibits significantly higher accuracy than other counterparts, which implies that MixNet-Copilot can find the most intensive pairs in all-to-all communication with high probability. Therefore, MixNet-Copilot offers an opportunity to proactively reconfigure the topology for the FP’s first all-to-all in advance.
@@ B.2. Details of Topology Reconfigurations
Figure 20 details the runtime reconfiguration methodology of MixNet. Each MoE layer involves four all-to-all communication phases. As discussed in §5.1, MixNet can deterministically characterize the traffic patterns for the second all-to-all communication in the forward pass (FP) and both all-to-all communications in the backward pass (BP).
Consequently, MixNet conceptually reconfigures the topology twice per MoE layer—once during FP and once during BP. However, for the first all-to-all communication in FP, MixNet cannot fully characterize the traffic matrix in advance due to the absence of runtime information at that stage of the iteration. To tackle this issue, MixNet performs an inaccurate reconfiguration based on partial estimates or reuses the topology from the previous layer. It then recalibrates the topology for the second all-to-all in FP with minor OCS adjustments, ensuring a more accurate configuration for subsequent communication phases.
@@ Appendix C ProtoType Details
As a result, we currently exclude this NIC activation time to calculate the actual training time in MixNet testbed experiments. Due to the limited number of GPUs, we cannot train the full MoE models as shown in Table 1. We only run 7 layers of Mixtral 8$\times$7B, 16 layers of LLaMA-MoE, and 12 layers of Qwen-MoE.
@@ Appendix D Large-Scale Simulation Details
D.1. Simulated MoE Models and Parallelization Strategies
@@ D.2. Cost Analysis Details
D.3. Different EPS Link Options
The OCS portion of MixNet requires optical transceivers with pluggable fibers to allow optical switching. For the EPS part of MixNet, especially short-reach rack-scale links between the servers and ToR switches, Direct Attach Copper (DAC) cables or Active Optical Cables (AOC) are more cost-effective alternatives to optical transceivers plus fibers (typically used for long-reach links). We analyze the cost implications of these link options in Figure 24. The results show that replacing the EPS links with DAC or AOC slightly reduce the costs for both fat-tree interconnect and MixNet. Most importantly, the cost effectiveness of MixNet is orthogonal to the choices of EPS links, and maintains significant cost advantages over fat-tree topology. For example, with 400 Gbps DAC cables option in a 4096-GPU cluster, MixNet achieves 2.2$\times$ lower total cost compared to fat-tree topology.
@@ D.4. Training Speed Ups of Mixtral Models with Larger Batch Sizes
We further evaluate MixNet’s performance with larger batch sizes using two Mixtral MoE models (Mixtral 8$\times$7B and Mixtral 8$\times$22B). For each model, we test batch sizes of 32 and 64 across varying network bandwidths (100-800 Gbps). As shown in Figure 25, MixNet consistently outperforms TopoOpt under all configurations. Specifically, MixNet achieves an average speedup of 1.8$\times$ for Mixtral-8x7B with a batch size of 32 and 2.0$\times$ with a batch size of 64, as training becomes more communication-intensive compared to the settings in Figure 12. Furthermore, we observe that as link bandwidth increases, MixNet’s performance gradually approaches that of Fat-tree and Rail-optimized architectures.
@@
Figure 26. [Simulation] Scalability analysis of MixNet with different cluster sizes.
Figure 27. [Simulation] Impact of optical degree $\alpha$ in MixNet.
@@ D.5. Scalability
We demonstrate the scalability of MixNet in Figure 26(a). The Mixtral 8$\times$7B model is evaluated at 400 Gbps bandwidth, with the cluster size varying from 128 servers to 4,096 servers, covering up to 32768 GPUs. MixNet demonstrates scalability by fundamentally relaxing the port limits of OCS through the design of several decentralized regionally reconfigurable domains, allowing it to scale similarly to a fat-tree topology. Our results show that MixNet scales effectively with increasing number of GPUs, achieving training throughput comparable to both non-blocking Fat-tree and Rail-optimized topologies in terms of tokens processed per second. We further present the performance-cost comparison in Figure 26(b), which shows that MixNet consistently achieves a superior performance-cost trade-off—approximately 2$\times$ higher performance-per-dollar—compared to Fat-tree and Rail-optimized topologies as the number of GPUs increases. This suggests that MixNet maintains the training cost-effectiveness even as the cluster size grows.
@@ D.6. Impact of Optical Degree
We show the impact of the optical degree on MixNet’s performance in Figure 27. We evaluate the Mixtral 8$\times$22B model on a cluster of 128 servers with 100 Gbps link bandwidth. The optical degree $\alpha$ in MixNet is varied to adjust its connectivity in the OCS. We reduce the bandwidth of each electronic port when increasing their number, to ensure a cost-equivalent comparison. Our findings show that, as the optical degree increases, MixNet further reduces iteration time, as more communication-intensive GPU pairs can be provisioned with dedicated high-bandwidth optical circuits.
D.7. Impact of Reconfiguration Latency
@@
To investigate MixNet’s sensitivity to OCS reconfiguration latency, we evaluate the Mixtral 8$\times$22B model on a cluster of 128 servers with 400 Gbps link bandwidth, varying the reconfiguration latency from 1 µs to 10 s. Figure 28 shows the normalized iteration time. MixNet assumes the use of a millisecond-scale reconfigurable OCS (25 ms) in its current implementation. We observe that further reducing reconfiguration latency does not yield significant performance gains, as the OCS reconfiguration process can already be fully hidden during the computation phase. However, provisioned with microsecond-scale reconfigurable OCS, MixNet can enable fully accurate topology reconfigurations for the first all-to-all communication in the forward pass (FP), resulting in marginal performance improvements in this specific phase. On the other hand, when the reconfiguration latency exceeds 1000 ms, performance degrades obviously, as the OCS reconfiguration may not be hidden and starts to block the training process. As a result, MixNet does not perform well with second-scale reconfigurable OCS systems.